OpenStudy (anonymous):

can some one give me a hand with this? 7^-x * (1/7)^(9x-2)= 49^(x^2-5)

OpenStudy (gorv):

\[\frac{ 1 }{ x }=x^{-1}\]

OpenStudy (gorv):

u know that ???

OpenStudy (anonymous):

i know that but i never thought i could solve it like that...............

OpenStudy (anonymous):

thank you!

OpenStudy (shinalcantara):

\[(7^{-x})(\frac{ 1 }{ 7 }) ^{9x-2}=49^{x^2-5}\] is that the given?

OpenStudy (anonymous):

yeah, i'm still stuck though

OpenStudy (anonymous):

how do u get id of the \[1^{9x-2}\]

OpenStudy (anonymous):

rid*

OpenStudy (shinalcantara):

first thing to do, express each term to the base of 7

OpenStudy (shinalcantara):

follow what gorv has pointed out above. express (1/7)^(9x-2) to a base of 7

OpenStudy (shinalcantara):

for the right side, 49 is a perfect square. express it also to the base of 7

OpenStudy (shinalcantara):

\[\frac{ 1 }{ 7 } \rightarrow 7^{-1}\]

OpenStudy (anonymous):

\[7^{-10x+2} = 7^{2x^2-10}\] is that right?

OpenStudy (shinalcantara):

yes it's correct